# The Early Bird Catches the Leak: Unveiling Timing Side Channels in

### v3

@@ Abstract
The wide deployment of Large Language Models (LLMs) has given rise to strong demands for optimizing their inference performance. Today’s techniques serving this purpose primarily focus on reducing latency and improving throughput through algorithmic and hardware enhancements, while largely overlooking their privacy side effects, particularly in a multi-user environment. In our research, for the first time, we discovered a set of new timing side channels in LLM systems, arising from shared caches and GPU memory allocations, which can be exploited to infer both confidential system prompts and those issued by other users. These vulnerabilities echo security challenges observed in traditional computing systems, highlighting an urgent need to address potential information leakage in LLM serving infrastructures.
In this paper, we report novel attack strategies designed to exploit such timing side channels inherent in LLM deployments, specifically targeting the Key-Value (KV) cache and semantic cache widely used to enhance LLM inference performance. Our approach leverages timing measurements and classification models to detect cache hits, allowing an adversary to infer private prompts with high accuracy. We also propose a token-by-token search algorithm to efficiently recover shared prompt prefixes in the caches, showing the feasibility of stealing system prompts and those produced by peer users. Our experimental studies on black-box testing of popular online LLM services demonstrate that such privacy risks are completely realistic, with significant consequences. Our findings underscore the need for robust mitigation to protect LLM systems against such emerging threats.
1 Introduction
@@
To address these challenges, we developed various attack strategies to exploit LLM side channels. Specifically, we use a classification model to detect token-level KV cache hits based on offline timing measurements. By dynamically adjusting the timing threshold at runtime, we observe that online detection accuracy can be substantially improved with only a few repeated trials. To reduce the search space for the KV cache channel, we propose an incremental search algorithm that capitalizes on the requirement for prompts to share a common prefix, allowing us to recover the victim’s prompt token by token. For the semantic cache channel, we design an algorithm to select the most representative prompts as the attacker’s requests, given a targeted semantic focus. To minimize interference from the attacker’s own requests, we introduce a mechanism to clear cached data via batches of irrelevant requests. For the semantic cache, our method also ensures the attacker’s requests remain distinct by computing their semantic similarities.
@@
For the prompt stealing attack, our evaluation indicates that the accuracy of detecting per-token cache hits or misses in the KV cache is 0.99, with a false positive rate (FPR) of 0.003. Using the incremental search algorithm, we recovered the system prompt token by token, requiring an average of 111.46 queries per recovered token. This approach achieved an average recovery accuracy of 89.0% and a corresponding FPR of 0.04. For the peeping neighbor attack, our measurements show an 81.4% accuracy in distinguishing hits from misses, with an average FPR of 0.045 in a single trial. This accuracy improved to 95.4% with a 0.056 FPR after 5 trials under GPTCache’s default settings. We further observed that it is possible to infer the documents processed by a victim user in a vulnerable LLM application, even when using standard commodity LLM services. Moreover, our black-box study of existing online services shows that popular LLM systems—such as Claude, DeepSeek, and Azure OpenAI—employ KV or semantic cache sharing to cut costs, rendering them susceptible to timing side-channel attacks.
Finally, we propose initial defenses against these side-channel risks. To mitigate KV cache leakage, we recommend sharing prefix caches only in batches of at least $k$ tokens ($k=2,3,4$, etc.). Although this increases the prompt search space and thus the required number of guesses, the larger timing differences for sharing multiple tokens also make classifiers more robust. Consequently, attacks remain accurate but incur higher query overhead. To address semantic cache leakage, we advise anonymizing privacy-related content in user inputs before performing semantic-similarity searches. Preliminary experiments show this measure adds modest overhead (around 4%).
@@
Responsible disclosure. We disclosed our findings to all relevant developers (SGLang, GPTCache, etc.) and LLM service providers (OpenAI, Claude, Google Gemini, etc.) upon identifying the side channels in September 2024. At the time of this manuscript’s preparation, we received positive responses from the SGLang team, which noted that we were among the first two groups to report this issue, both within the same week. Moreover, we were the first to raise the topic during the SGLang development meeting, and we are now working closely with their team on a resolution. We will make all the code and datasets necessary to reproduce our experiments publicly available once the paper is published.
2 Background
2.1 LLM Serving Systems
@@ 2.3 Serving Backend
2.4 Threat Model
In this paper, we examine the security implications of deploying a shared LLM to serve multiple users or applications within a single computing system. Specifically, we consider two main scenarios. First, an LLM service provider offers public APIs that registered users can employ to send requests, all of which are processed by the same underlying serving system. In this context, a victim user may establish a proprietary system prompt to power a widely used LLM application. Meanwhile, an attacker could leverage the same LLM APIs to infer this system prompt, thereby gaining potential financial benefits or circumventing the safety instructions encoded in the prompt. Second, public LLM applications—such as chatbots (e.g., OpenAI’s GPT-4) or document analysis services (e.g., AnythingLLM [7], Klu [15], etc.)—handle concurrent requests from multiple users via the same LLM serving system. If the application itself relies on a public LLM API, these requests are generally routed through the same developer’s API key. An attacker could register as a user of such an application to discover whether specific requests have been submitted by others. For instance, they might seek to determine whether a user has shown interest in a particular topic or uploaded a specific file. They could also monitor requests over time to detect private attributes or sensitive personally identifiable information (PII) (Table 2). In both scenarios, the attacker’s and the victim’s requests share the same platform and thus make use of the same caches. This shared environment can produce interference and create observable timing side channels—the core subject of our investigation. The attacker needs only black-box access to the underlying model, without knowledge of its architecture or weight parameters. However, the attacker must first examine the system’s leakage profile in an offline phase, analyzing how different inputs affect timing. This analysis helps them craft inputs that exploit the timing discrepancies introduced by cache sharing.
@@ 3.1 Overview
3.2 Prompt Stealing Attacks (PSA)
@@
prefill time (miss) =#$=\#tokens\times\frac{\#parameters}{\text{GPU compute bandwidth}}$
@@
prefill time (hit) =#$=\#tokens\times\frac{\text{KV cache per token}}{\text{GPU memory bandwidth}}$
@@
These gaps become larger for more complex models or when serving multiple requests concurrently. For instance, on a Llama-2-70B-Chat-GPTQ model (70 billion parameters at 4-bit precision), the prefill time for a single token miss is about 0.45 ms, while a hit is roughly 0.22 $\mu$s. These timing differences underpin the prompt stealing attack, which leverages KV cache sharing. Specifically, an attacker sends a request to the LLM and observes TTFT to detect whether the victim’s prefixes match. Since KV cache sharing occurs only for prompts with the same prefix, we devised an incremental search algorithm to recover prompts on a token-by-token basis. We present this algorithm and assess its real-world performance in Section 4.1 and Section 4.3.
3.3 Peeping neighbor Attacks (PNA)
@@
The side channel. Unlike KV caching, which reuses data only for identical prompt prefixes, semantic caching allows reuse based on semantic similarity beyond a predefined threshold. However, sharing a semantic cache among multiple users can inadvertently reveal their requests. Cache hits provide responses in mere milliseconds, whereas cache misses can take several seconds—creating a clear timing difference that can be exploited by an attacker. This discrepancy enables the attacker to infer the semantics of concurrent requests issued by nearby users, a scenario we refer to as the peeping neighbor attack. Despite this, when the attacker tries to match a victim’s request semantically, the attacker’s own requests may also be cached, introducing noise into subsequent attempts. To address this challenge, we propose an efficient search algorithm that both minimizes the caching effects of the attacker’s own requests and improves the detection rate for the victim’s request. We describe this algorithm and illustrate how it can recover private information from neighboring users’ prompts in Section 4.2.
4 Side-channel Analysis and Evaluation
@@
In this section, we present our empirical analysis of the identified side channels and describe strategies for their efficient exploitation. All experiments were conducted on a Lenovo server equipped with two Intel Xeon E5-2678 v3 CPUs (12 cores at 2.50 GHz each), 100 GB DDR4 memory, and two NVIDIA A100-PCIE GPUs (40 GB memory each). The system ran Ubuntu 22.04 (kernel 5.15.0-125-generic) with GPU driver 550.127.05, CUDA 12.4, and PyTorch 2.4.0. We used open-source models from the Llama family as the underlying LLM, adhering to their default hardware and software configurations for all evaluations.
4.1 Analysis on PSA
@@
Attack overview. With increasing concerns about client data leakage in public LLM services, enterprise LLMs have become increasingly popular [19, 22]. In our study, we focus on a use case where the LLM API service is constructed from open-source projects within a local network environment. As described in Section 2.4, we consider a scenario in which a victim develops a popular LLM application (e.g., a chatbot) using a proprietary system prompt via the LLM service. The attacker interacts with this LLM application over the local network, measures the TTFT, and attempts to uncover the system prompt based on timing discrepancies. Specifically, the LLM service uses the SGLang backend API server [30], which supports KV cache sharing for common prefixes. Notably, LLM API servers often allow users to define various roles, such as “system” and “user”, within their requests. The victim’s LLM chatbot is built on the FastChat framework [92]. As shown in Figure 1, FastChat supports both direct and synthesized operation modes. In direct mode, the user sends requests directly to the SGLang backend, whereas in synthesized mode, the full prompt is created by concatenating messages from each role according to predefined templates.
@@
Figure 1: LLM chatbots like FastChat allow user input through both direct requests (top) and synthesized requests via a template (bottom).
@@
(a) LLaMA-3.1-8B-Instruct
(b) LLaMA-2-70B-GPTQ
@@
Figure 3: Latency distribution of one token hit and miss. We used 2 representative models of different sizes: llama3.1-8B-Instruct and llama2-70B-GPTQ. We use the llama2-70B-GPTQ model for subsequent evaluations.
@@
Characterizing the leakage. We began by examining the timing difference between a cache hit and a miss for a single token, comparing multiple model sizes. Specifically, we tested SGLang v0.3.0 [30], which organizes KV cache blocks in a radix tree to facilitate efficient prefix matching and reuse. Two models were evaluated: llama3.1-8B-Instruct and llama2-70B-GPTQ. System prompts of varying lengths were derived from the gabrielchua/system-prompt-leakage [32] dataset on Hugging Face [32]. We measured TTFTs under conditions where the shared-prefix token count differed by exactly one—signifying cache hits versus misses—across 4,000 runs. Figure 3 illustrates the resulting time distributions, revealing a pronounced distinction between hit and miss scenarios for both models.
Based on these observations, a straightforward classifier can be built to categorize a token as a hit if its latency is below a certain threshold. As prompts grow longer, the latency also tends to rise due to increased computational demands. Consequently, the threshold must be adjusted according to the token’s position in the prompt, as well as the particular hardware and model used.
In real-world evaluations of our classifier, we noted that TTFT often varies due to factors such as GPU system noise and fluctuations in voltage and power, weakening the effectiveness of a fixed classification threshold. To address this challenge, we introduce a dynamic calibration scheme that continuously adjusts the threshold in real time, enhancing the classifier’s robustness. The primary mechanism involves simultaneously collecting TTFT data from known requests (i.e., requests without appended predicted tokens) within a brief time window alongside the targeted request’s TTFT. These concurrent measurements establish a baseline threshold representing real-time system performance. Based on this baseline, the threshold is dynamically updated at runtime to boost accuracy.
@@
To illustrate the classifier’s effectiveness, consider an example using the LLaMA-2-70B-GPTQ model under our evaluation settings. We build a classifier to determine whether the 10-th token hits the KV cache. Using the profiled timing distribution, we start with an initial threshold of 0.07971 seconds and then dynamically refine it based on the system’s real-time performance. We test the classifier using 4,000 hits and 4,000 misses drawn from randomly selected system prompts, achieving a TPR of 0.88 and a FPR of 0.10. To further mitigate noise, we employ multi-sampling by collecting $n$ TTFT samples per token. The token is classified as a hit only if the number of hit detections exceeds a threshold $k$. With $n=10$ and $k=5$, the final classifier attains a TPR of 0.99 and an FPR of 0.003.
End-to-end attacks. We built a local chatbot using the FastChat framework as the victim application. Its backend API server is configured with SGLang v0.3.0, which supports KV cache sharing for common prefixes. In this evaluation, we used the gabrielchua/system-prompt-leakage [32] dataset from Hugging Face [32], containing over 350,000 synthetic system prompts. Following the setup in Section 2.4, we assume the attacker has a public dataset of system prompts that mirrors those of the victim. Therefore, we randomly selected 200 prompts from the dataset as the victim’s prompts and used the remainder for the attacker.
Figure 4: Efficient token-by-token request recovery.
@@
To streamline the search process, as illustrated in Figure 4, we propose an incremental token-by-token approach to recover the target prompt. This approach relies on multiple components: a series of token classifiers for validation, a next-token predictor to estimate token probabilities, and a sampler that selects candidate tokens based on the predictor’s temperature setting. The next-token predictor is fine-tuned on the public system prompt dataset available to the attacker, allowing it to forecast the next token given the already retrieved tokens. For each candidate token at position $i$, we feed the partially reconstructed prompt into the LLM to obtain TTFT data, which then serves as the input to the corresponding $classifier_{i}$. If the classifier identifies this token as a cache hit, the token is appended to the prompt; otherwise, a repetition penalty is applied by adjusting the probability distribution of the current token (in our implementation, this penalty is applied by halving the sampling probability for an incorrect token in the next round).
@@
In constructing the next-token predictor, we adopt LLaMA-3.1-8B-Instruct [59] as the base model and fine-tune it using the attacker’s training dataset. Given the partially recovered prompt and the chosen temperature, the predictor generates a probability distribution for the next token. Temperature scaling introduces variability into the prediction. Our predictor is not heavily optimized; it was fine-tuned on a single NVIDIA A100-PCIE GPU for only a few hours with limited training resources. With larger models, more epochs, and bigger batch sizes, the predictor’s performance could be further improved.
@@
$\bullet$ Flushing the caches through eviction. We observed SGLang provides a $flush\_cache$ API [47] that efficiently clears the cache. However, for our end-to-end attack scenario, we chose not to use this API, as it is unlikely to be accessible to attackers in real-world environments. Instead, we employed a more robust method of evicting the KV cache by issuing batches of irrelevant requests. Under default SGLang settings, sending 15 such requests (each containing about 300 tokens) was sufficient to trigger eviction in about 5 seconds. This approach proved successful in 100% of our 1,000 tests.
@@
response = client.chat.completions.create(...)
@@
if data.get("end_of_sequence", False):
@@
Although the FPR for detecting a single token’s cache hit or miss is low (0.003), the FPR can accumulate if the next-token predictor repeatedly suggests incorrect tokens, potentially leading to an erroneous token recovery. To address this, we introduce a cross-verification step for each candidate token. Specifically, we send another query that omits the token to estimate the hit-based TTFT. Next, we compare the TTFT of this query with that of the predicted prompt. If the TTFT difference exceeds a preset threshold corresponding to one token’s prefill time, the current prediction is deemed incorrect, and we proceed to the next guess. This token recovery process continues until either a predetermined number of tokens is successfully retrieved or the maximum allotted attack queries is reached.
We evaluated the recovery accuracy and the average number of queries needed to retrieve each token. Our results show a success rate of 89.0%, an FPR of 0.04, and an average of 5.57 guesses with 111.46 attack queries per recovered token. Out of 200 victim prompts, we successfully recovered an average of 11.58 tokens for the top 100 prompts, and 17.9 tokens on average for the top 50 prompts. The maximum number of tokens recovered for a single prompt was 81, achieved with just 513 total guesses. Table 1 presents several examples of target system prompts alongside the prompts recovered via PSA. In real-world attacks, the attacker can recover additional tokens by deploying a more advanced next-token predictor and increasing the maximum number of attack queries. A demo for the end-to-end attack is presented in our website [37].
Table 1: Examples of recovered system prompts, including the number of attack queries and recovered tokens. We only listed travel planning-related prompts to demonstrate the PSA’s ability to recover diverse expressions.

### v4

@@ Abstract
The wide deployment of Large Language Models (LLMs) has given rise to strong demands for optimizing their inference performance. Today’s techniques serving this purpose primarily focus on reducing latency and improving throughput through algorithmic and hardware enhancements, while largely overlooking their privacy side effects, particularly in a multi-user environment. In our research, for the first time, we discovered a set of new timing side channels in LLM systems, arising from shared caches and GPU memory allocations, which can be exploited to infer both confidential system prompts and those issued by other users. These vulnerabilities echo security challenges observed in traditional computing systems, highlighting an urgent need to address potential information leakage in LLM serving infrastructures. In this paper, we report novel attack strategies designed to exploit such timing side channels inherent in LLM deployments, specifically targeting the Key-Value (KV) cache and semantic cache widely used to enhance LLM inference performance. Our approach leverages timing measurements and classification models to detect cache hits, allowing an adversary to infer private prompts with high accuracy. We also propose a token-by-token search algorithm to efficiently recover shared prompt prefixes in the caches, showing the feasibility of stealing system prompts and those produced by peer users. Our experimental studies on black-box testing of popular online LLM services demonstrate that such privacy risks are completely realistic, with significant consequences. Our findings underscore the need for robust mitigation to protect LLM systems against such emerging threats.
Index Terms:
LLM, KV cache, Semantic cache, Side channels
I Introduction
@@
To address these challenges, we developed various attack strategies to exploit LLM side channels. Specifically, we use a threshold-based classification model to detect token-level KV cache hits based on offline timing measurements. We observe that online detection accuracy can be substantially improved with only a few repeated trials. To reduce the search space for the KV cache channel, we propose an incremental search algorithm that capitalizes on the requirement for prompts to share a common prefix, allowing us to recover the victim’s prompt token by token. For the semantic cache channel, we design an algorithm to select the most representative prompts as the attacker’s requests, given a targeted semantic focus. To minimize interference from the attacker’s own requests, we introduce a mechanism to clear cached data via batches of irrelevant requests. For the semantic cache, our method also ensures the attacker’s requests remain distinct by computing their semantic similarities.
@@
For the prompt stealing attack, our evaluation indicates that the accuracy of detecting per-token cache hits or misses in the KV cache is 99%, with a false positive rate (FPR) of 0.003. Using the incremental search algorithm, we recovered the system prompt token by token, requiring an average of 111.46 queries per recovered token. This approach achieved an average recovery accuracy of 89.0% and a corresponding FPR of 0.04. For the peeping neighbor attack, our measurements show an 81.4% accuracy in distinguishing hits from misses, with an average FPR of 0.045 in a single trial. This accuracy improved to 95.4% with a 0.056 FPR after 5 trials under GPTCache’s default settings. We further observed that it is possible to infer the documents processed by a victim user in a vulnerable LLM application, even when using standard commodity LLM services. Moreover, our black-box study of existing online services shows that popular LLM systems—such as Claude, DeepSeek, and Azure OpenAI—employ KV or semantic cache sharing to cut costs, rendering them susceptible to timing side-channel attacks.
Finally, we propose initial defenses against these side-channel risks. To mitigate KV cache leakage, we recommend sharing prefix caches only in batches of at least $k$ tokens ($k=2,3,4$, etc.). Although this increases the prompt search space and thus the required number of guesses, the larger timing differences for sharing multiple tokens also make classifiers more robust. Consequently, attacks remain accurate but incur higher query overhead. To address semantic cache leakage, we advise anonymizing privacy-related content in user inputs before performing semantic-similarity searches. Preliminary experiments show that this measure adds modest overhead (around 4%).
@@
Responsible disclosure. We disclosed our findings to all relevant developers (SGLang, GPTCache, etc.) and LLM service providers (OpenAI, Claude, Google Gemini, etc.) upon identifying the side channels in September 2024. At the time of this manuscript’s preparation, we received positive responses from the SGLang team, which noted that we were among the first two groups to report this issue, both within the same week. Moreover, we were the first to raise the topic during the SGLang development meeting, and we are now working closely with their team on a resolution.
TABLE I: Comparisons with closely-related works.
Independent of scheduling policy? Independent of prompt templates? Cache eviction supported? Real-world measurement? KV cache channel? Semantic cache channel? Mitigations
PROMPTLEAK [29]
InputSnatch [30]
This work
Comparison with concurrent and follow-up works (Table I). Concurrently and independently to our research, Wu et al. proposed PROMPTLEAK [29], an attack that exploits the Longest Prefix Match (LPM) scheduling policy in SGLang, which prioritizes requests with longer prefix matches, to leak user prompts. Their method performs collision attacks by sending carefully crafted batched prompts. When a request shares more KV cache entries than others, SGLang prioritizes it, enabling adversaries to extract victim tokens. In contrast, our work identifies a more general KV cache side channel that does not depend on LPM-based scheduling and thus remains effective even without prefix-based prioritization. Moreover, we are the first to reveal a semantic cache side channel, propose practical mitigations, and conduct real-world measurements to assess leakage risks in commercial LLMs.
Inspired by our findings, Gu et al. conducted a subsequent large-scale measurement study on prompt caching in real-world LLM services [31], identifying its presence in 8 out of 17 evaluated providers.
More recently, Zheng et al. investigated timing side channels in LLMs through the InputSnatch attack [30]. However, their work lacks the optimized search strategy we introduce for efficient request recovery and suffers from practical limitations. Specifically, their KV cache attacks on vLLM can extract blocks of 16 tokens but are limited to template-dependent scenarios, requiring attackers to have prior knowledge of fixed input structures. Additionally, their semantic cache attacks on GPTCache aim to infer patterns from retrieved documents but do not address self-induced interference from the attackers’ own queries and lack eviction strategies to mitigate adversarial noise. In contrast, our approach enables template-free extraction, independent of application-specific formats, and incorporates eviction-controlled probing to systematically eliminate interference. Furthermore, we validate the effectiveness of our method in a realistic setting: a document summarization service powered by a commercial LLM API.
Availability. All the code and datasets necessary to reproduce our experiments are publicly available at: https://github.com/Maxppddcsz/llm-sidechannel. The demos for our attacks are available at: https://sites.google.com/view/early-bird-catches-the-leak/.
II Background
II-A LLM Serving Systems
@@ II-C Serving Backend
II-D Threat Model
In this paper, we examine the security implications of deploying a shared LLM to serve multiple users or applications within a single computing system. Specifically, we consider two main scenarios. First, an LLM service provider offers public APIs that registered users can employ to send requests, all of which are processed by the same underlying serving system. In this context, a victim user may establish a proprietary system prompt to power a widely used LLM application. Meanwhile, an attacker could leverage the same LLM APIs to infer this system prompt, thereby gaining potential financial benefits or circumventing the safety instructions encoded in the prompt. Second, public LLM applications—such as chatbots (e.g., OpenAI’s GPT-4) or document analysis services (e.g., AnythingLLM [39], Klu [40], etc.)—handle concurrent requests from multiple users via the same LLM serving system. If the application itself relies on a public LLM API, these requests are generally routed through the same developer’s API key. An attacker could register as a user of such an application to discover whether specific requests have been submitted by others. For instance, they might seek to determine whether a user has shown interest in a particular topic or uploaded a specific file. They could also monitor requests over time to detect private attributes or sensitive personally identifiable information (PII) (Table III). In both scenarios, the attacker’s and the victim’s requests share the same platform and thus make use of the same caches. This shared environment can produce interference and create observable timing side channels—the core subject of our investigation. The attacker needs only black-box access to the underlying model, without knowledge of its architecture or weight parameters. However, the attacker must first examine the system’s leakage profile in an offline phase, analyzing how different inputs affect timing. This analysis helps them craft inputs that exploit the timing discrepancies introduced by cache sharing.
@@ III-A Overview
III-B Prompt Stealing Attacks (PSA)
@@
prefill time (miss) $=\#tokens\times\frac{\#parameters}{\text{GPU compute bandwidth}}$
@@
prefill time (hit) $=\#tokens\times\frac{\text{KV cache per token}}{\text{GPU memory bandwidth}}$
@@
[…]
These timing differences underpin the prompt stealing attack, which leverages KV cache sharing. Specifically, an attacker sends a request to the LLM and observes TTFT to detect whether the victim’s prefixes match. Since KV cache sharing occurs only for prompts with the same prefix, we devised an incremental search algorithm to recover prompts on a token-by-token basis. We present this algorithm and evaluate its real-world efficiency in Section IV-A.
III-C Peeping Neighbor Attacks (PNA)
@@
The side channel. Unlike KV caching, which reuses data only for identical prompt prefixes, semantic caching allows reuse based on semantic similarity beyond a predefined threshold. However, sharing a semantic cache among multiple users can inadvertently reveal their requests. Cache hits provide responses in mere milliseconds, whereas cache misses can take several seconds—creating a clear timing difference that can be exploited by an attacker. This discrepancy enables the attacker to infer the semantics of concurrent requests issued by nearby users, a scenario we refer to as the peeping neighbor attack. Despite this, when the attacker tries to match a victim’s request semantically, the attacker’s own requests may also be cached, introducing noise into subsequent attempts. To address this challenge, we propose an efficient search algorithm that both minimizes the caching effects of the attacker’s own requests and improves the detection rate for the victim’s request. We describe this algorithm and illustrate how it can recover private information from neighboring users’ prompts in Section IV-B.
IV Side-channel Analysis and Evaluation
@@
In this section, we present our empirical analysis of the identified side channels and describe strategies for their efficient exploitation. All experiments were conducted on a Lenovo server equipped with two Intel Xeon E5-2678 v3 CPUs (12 cores at 2.50 GHz each), 100 GB DDR4 memory, and two NVIDIA A100 PCIE GPUs (80 GB memory each). The system ran Ubuntu 22.04 (kernel 5.15.0-125-generic) with GPU driver 550.127.05, CUDA 12.4, and PyTorch 2.4.0. We used open-source models from the Llama family as the underlying LLM, adhering to their default hardware and software configurations for all evaluations.
IV-A Analysis on PSA
@@
Attack setup. With increasing concerns about client data leakage in public LLM services, enterprise LLMs have become increasingly popular [52, 53]. In our study, we focus on a use case where the LLM API service is constructed from open-source projects within a local network environment. As described in Section II-D, we consider a scenario in which a victim develops a popular LLM application (e.g., a chatbot) using a proprietary system prompt via the LLM service. The attacker interacts with this LLM application over the local network, measures the TTFT, and attempts to uncover the system prompt based on timing discrepancies. Specifically, the LLM service uses the SGLang backend API server [26], which supports KV cache sharing for common prefixes. Notably, LLM API servers including OpenAI allow users to define various roles, such as “system” and “user”, within their requests [54], or send requests directly. Based on this capability, we categorize requests into two modes: synthesized and direct. The victim’s LLM chatbot, built on the FastChat framework [55], explicitly supports both, as shown in Figure 1. In direct mode, the user sends requests directly to the SGLang backend, whereas in synthesized mode, the full prompt is created by concatenating messages from each role according to predefined templates.
@@
Figure 1: LLM API servers like OpenAI allow user input through both direct requests (top) and synthesized requests via a template (bottom).
@@
(a) (a). Llama-3.1-8B-Instruct
(b) (b). Llama-3.1-70B-Instruct-GPTQ-INT4
@@
Figure 3: Latency distribution of one token hit and miss. We used 2 representative models of different sizes: Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct-GPTQ-INT4. We used the Llama-3.1-70B-Instruct-GPTQ-INT4 model for subsequent evaluations.
@@
Characterizing the leakage. We began by examining the timing difference between a cache hit and a miss for a single token, comparing multiple model sizes. Specifically, we tested SGLang v0.3.0 [26], which organizes KV cache blocks in a radix tree to facilitate efficient prefix matching and reuse. Two models were evaluated: Llama-3.1-8B-Instruct [57] and Llama-3.1-70B-Instruct-GPTQ-INT4 [50]. System prompts of varying lengths were derived from the gabrielchua/system-prompt-leakage [58] dataset on Hugging Face. We measured TTFTs using 15-token prompts where only the last token differed, resulting in a shared-prefix length differing by exactly one—signifying cache hits versus misses—across 4,000 runs. Figure 3 illustrates the resulting time distributions, revealing a pronounced distinction between hit and miss scenarios for both models.
Based on these observations, a straightforward classifier can be built to categorize the last token in the prompt as a hit if its latency is below a predefined threshold. However, as prompts grow longer, the latency also tends to rise due to increased computational demands, necessitating a distinct length-dependent threshold for each prompt length. To overcome this limitation, we measured the relative latency difference between hit and miss cases for last tokens across prompts of varying lengths (1–200 tokens). Our findings reveal that this difference is stable, independent of prompt length, and significantly larger than the variance within miss cases. This stability allows us to use a classifier with a single threshold instead of separate classifiers for different lengths, ensuring reliable hit detection.
In real-world evaluations of our classifier, we noted that TTFT varies due to factors such as GPU system noise and power fluctuations, weakening the effectiveness of a fixed classification threshold, as shown in Figure 3. To mitigate this, we first construct a miss prompt by appending a rare token to a known prompt (i.e., requests without predicted tokens). Then, within a brief time window, we simultaneously collect TTFT data for both the miss and target prompts within a brief time window, using their difference for threshold-based classification.
@@
To illustrate the classifier’s effectiveness, consider an example using the Llama-3.1-70B-Instruct-GPTQ-INT4 [50] model under our evaluation settings. We build a classifier to determine whether the last token of prompts hits the KV cache. We test the classifier using 4,000 hits and 4,000 misses drawn from randomly selected system prompts, achieving a TPR of 0.88 and a FPR of 0.10. To further mitigate noise, we employ multi-sampling by collecting $n$ TTFT samples per token. The token is classified as a hit only if the number of hit detections exceeds a threshold $k$. With $n=10$ and $k=5$, the final classifier attains a TPR of 0.99 and an FPR of 0.003.
End-to-end attacks. We built a local chatbot using the FastChat framework as the victim application. Its backend API server is configured with SGLang v0.3.0, which supports KV cache sharing for common prefixes. In this evaluation, we used the gabrielchua/system-prompt-leakage [58] dataset from Hugging Face, containing over 350,000 synthetic system prompts. We applied Llama-3.1-8B-Instruct [57] to compute embeddings for each prompt and visualized the results using Uniform Manifold Approximation and Projection (UMAP), which reduces dimensions of datasets and enables effective visualization of data distribution [59]. UMAP visualization confirmed substantial semantic diversity across the dataset, reflecting the varied topics and styles found in real-world prompt engineering. As shown in Figure 4, this semantic heterogeneity increases as the dataset expands, significantly complicating predictive modeling [60]. Consistent with the threat model detailed in Section II-D, we assume the attacker has a public dataset of system prompts that mirrors those of the victim. To more accurately reflect real-world conditions—-where production systems typically implement a small number of carefully crafted system prompts—-we randomly select only 200 prompts as the victim’s configuration while reserving the remainder as the attacker’s training corpus. This experimental design deliberately simulates the “needle in a haystack” challenge faced in realistic attack scenarios, where adversaries must identify specific target prompts within a substantially larger heterogeneous corpus.
(a) (a). Random sample of 10k prompts from the dataset
(b) (b). Complete dataset (over 350k prompts)
Figure 4: UMAP projection demonstrating the dataset’s heterogeneity: As sampling density increases, the visualization reveals expanded dimensional ranges with distinct semantic clusters and isolated points, highlighting the complex multidimensional structure of the prompt dataset.
@@
To streamline the search process, as illustrated in Figure 6, we propose an incremental token-by-token approach to recover the target prompt. This approach relies on multiple components: a classifier for validation, a next-token predictor to estimate token probabilities, and a sampler that selects candidate tokens based on the predictor’s temperature setting. The next-token predictor is fine-tuned on the public system prompt dataset available to the attacker, allowing it to forecast the next token given the already retrieved tokens. For each candidate token at position $i$, we feed the partially reconstructed prompt and miss prompt into the LLM to obtain TTFT difference (same method in Section IV-A), which then serves as the input to the classifierclassifier. If the classifier identifies this token as a cache hit, the token is appended to the prompt; otherwise, a repetition penalty is applied by adjusting the probability distribution of the current token (in our implementation, this penalty is applied by halving the sampling probability for an incorrect token in the next round).
@@
In constructing the next-token predictor, we adopt Llama-3.1-8B-Instruct [57] as base model, chosen as it shares an identical tokenizer with the victim LLM. We fine-tune this model using the attacker’s training dataset. Given the partially recovered prompt and the chosen temperature, the predictor generates a probability distribution for the next token. Temperature scaling introduces variability into the prediction. Our predictor is not heavily optimized; it was fine-tuned on a single NVIDIA A100 PCIE GPU for only a few hours with limited training resources. With larger models, more epochs, and bigger batch sizes, the predictor’s performance could be further improved.
@@
response = client.chat.completions
.create(...)
@@
if data.get(
"end_of_sequence",
False):
@@
$\bullet$ Flushing the caches through eviction. We observed SGLang provides a flush_cacheflush\_cache API [61] that efficiently clears the cache. However, for our end-to-end attack scenario, we chose not to use this API, as it is unlikely to be accessible to attackers in real-world environments. Instead, we employed a more robust method of evicting the KV cache by issuing batches of irrelevant requests. Under default SGLang settings, sending 15 such requests (each containing about 200 tokens) was sufficient to trigger eviction in about 5 seconds. This approach proved successful in 100% of our 10,000 tests.
Figure 6: Efficient token-by-token request recovery.
We evaluated both the token recovery accuracy and the average number of queries required per token. The results indicate a success rate of 89.0%, an FPR of 0.04, and an average of 5.57 guesses with 111.46 attack queries needed per recovered token when $n=10$. The corresponding cost per token recovery is approximately \$0.16 using OpenAI o1, \$0.012 with OpenAI o3-mini, and \$0.001 with Deepseek-Chat (as of March 5, 2025). Out of 200 victim prompts, we successfully recovered an average of 11.58 tokens for the top 100 prompts, and 17.9 tokens on average for the top 50 prompts. The maximum number of tokens recovered for a single prompt was 81, achieved with just 513 total guesses. Table II presents several examples of target system prompts alongside the prompts recovered via PSA. Notably, we limited the token prediction attempts to a maximum of 80 per position, ceasing further attempts if the correct token was not identified within these trials. Consequently, the primary constraint in reconstructing entire system prompts stems from the precision of the predictor rather than the accuracy of the classifier—an aspect that falls outside the scope of this study. In real-world attacks, the attacker can recover additional tokens by deploying a more advanced next-token predictor and increasing the maximum number of attack queries. A demo for the end-to-end attack is presented in our website [62].
TABLE II: Examples of recovered system prompts, including the number of attack queries and recovered tokens. We only listed travel planning-related prompts to demonstrate the PSA’s ability to recover diverse expressions.